The width of a rectangle is 10 meters and its area is 120 square meters. What is the length of the diagonal?
Answer by checkley77(12844)

A=LW
120=L*10
L=120/10
L=12 M. ANS. FOR THE OTHER SIDE.
DIAGONAL IS:
10^2+12^2=C^2
100+144=C^2
C^2=244
C=SQRT244
C=15.62 M. ANS. FOR THE DIAGONAL.